OpenStudy (anonymous):

how do i solve the inverse tangent of 6?

OpenStudy (psymon):

I think this problem actually requires that you have the rest of the problem written. I had an idea on how to solveit, but it requires the entire problem you had xD

OpenStudy (anonymous):

lol it looks less scary this way. But Here's the entire equation folks: \[Cos (\frac{ 1 }{ 2 } \tan^{-1} 6)\]

OpenStudy (psymon):

Well, because when you have a problem like this, its actual triangles and not just calculator work. This is a triangle problem, so thats why you dont need a calculator forit. If it were just tan^(-1)6 then its all calculator.

OpenStudy (psymon):

Okay, so we need to say that the whole (1/2)tan^(-1)6 is going to be θ. So we will, eventually, have cos(θ) = ?. So now how to deal with the 1/2arctan bleck. So we need a triangle for this. First, though, if arctan (sorry, I like writing arc better), so if arctan6 = θ, then tanθ = 6. That being said, in terms of a triangle, tangent is opposite over adjacent. So inside of the parenthesis, I have (1/2)tanθ = 6. Solving for tanθ, I have tanθ = 12. Now HERE is where the triangle comes in. *draws*

OpenStudy (psymon):

 What we basically have done is said that the cosine of a triangle whose tangent is 6, has an angle of 1/√(145), or basically √(145)/145.

OpenStudy (psymon):

Could be wrong, but I have an idea what to do with it x_x

OpenStudy (psymon):

Wait, wait, wait, wait, I got it!!!

OpenStudy (anonymous):

@Psymon thank you so much for that explanation. This actually makes sense- my only concern is that i know i need to use a sum or difference formula for this kind of problem because there is a MUCH simpler problem in the book. This is that problem on steroids *_*

OpenStudy (psymon):

I need to correct myself, I know how to solve this. Can I re-explain? x_x

OpenStudy (anonymous):

go ahead! :)

OpenStudy (psymon):

Okay, we need to do a half-angle formula! We need tosay that θ is arctan(6), but we needour cos(θ/2) formula :P

OpenStudy (anonymous):

HOW DID I NOT SEE THAT. Man!!

OpenStudy (psymon):

I dont know, not sure how I didnt, haha.

OpenStudy (anonymous):

ok ill plug it into the cos\[\cos \theta \] formula and check my answer with you if you don't mind :) I make silly calculator mistakes

OpenStudy (psymon):

Haha, okay. But yeah, I just checked with my calculator. I plugged in the triangle method along with the flat out put it inthe calculator method and I got them both to match :P

OpenStudy (anonymous):

did you use the formula \[\sqrt{\frac{ 1+\cos \alpha }{ 2 }}\] ?

OpenStudy (anonymous):

i used this one because 6 is in quadrant 4 :p

OpenStudy (psymon):

But cos is positive in quadrant 4 :P

OpenStudy (anonymous):

why wouldn't it be positive? O_0

OpenStudy (psymon):

Sorry, I read it wrong :P But yeah, youre right, haha. But its √[(1+cos2θ)/2]

OpenStudy (psymon):

Wait....

OpenStudy (psymon):

Oops, youre rightxD

OpenStudy (psymon):

Got it confused with power-reducing

OpenStudy (anonymous):

haha that's ok, so my theta in the half angle formula is arctan 6??

OpenStudy (psymon):

Yes, but from there we need to get the cos of that, which is done with the triangle :P

OpenStudy (psymon):